angle A and angle B are complimentary. Find m of angle A and m of angle B.
m of angle A = 7x+1
m of angle B = 5x-7

Answer by miss_elaine17(1)

First, complimentary means the two angles equal 90 degrees.
Next, you have to solve for x.
7x+1+5x-7=90
(The 90 came from complimentary, which, again, means 90 degrees)
Combine like-terms, which are the numbers before the equals sign.
So, now your equation looks like:
12x-6=90
Add the 6 to all numbers without an x.
12x-6=90
+6 +6
12x =96
You do this to cancel out the extra 6.
Next, divide 12x and 96 by 12.
12x/12 =96/12
You get the 'divided by 12' from the number with the x. That goes for any time you complete an equation such as this one. (Example: if it would have been 6x =40, you would've divided 6x and 40 by 6. If it would have been 5y =50, you would've divided 5x and 50 by 5)
Okay, since 12 divided by 12 is 0, you don't write 12, just bring down your x.
96/12=8
You should be at this step by now:
x=8
If you got 8, then congratulations, you solved it!

If you would like to check your work, that is always an option.
Do so like this:
Go back to your original problem.
(7x+1 and 5x-7)
remember that x=8
7(8)+1=57
5(8)-7=33
57+33=90
Hope this helped!